OpenStudy (anonymous):

Find the sum of the measures of the interior angles of a regular 16-gon. Answer 2520 2848 2880 3240

OpenStudy (anonymous):

2520

OpenStudy (anonymous):

(180-360/16)*16=2520

OpenStudy (anonymous):

Exterior = 360/16 =22.50 so interior 157.50 *16 = 2520